# LaF-GRPO: In-Situ Navigation Instruction Generation for the Visually Impaired via GRPO with LLM-as-Follower Reward

### v1

@@ 4 Benchmark: NIG4VI
R2R (Anderson et al., 2018) High 21k
REVERIE (Qi et al., 2020) High 10k / 6k
UrbanWalk (Huang et al., 2022) Detailed 2.6k
Merchant et al. (2024) Detailed 48
VIALM (Zhao et al., 2024) Detailed 200
WAD (Yuan et al., 2025) Detailed 12k / 120k
@@
Table 1: Comparison of the NIG4VI dataset with existing benchmarks. Unlike WAD, NIG4VI employs geospatial coordinates for high spatial accuracy evaluation, and surpasses UrbanWalk through open-ended instruction generation that benchmarks VLMs’ capacity to produce natural navigation guidance.
Figure 3: DeepSeek-R1 refines initial predictions from GPT-4o, Claude-3.5, and Gemini-2 into coherent instructions, using Vision-R1’s modality bridging method (Huang et al., 2025). These are then reviewed and modified by human annotators for quality and accuracy.
@@
We introduce the NIG4VI benchmark to address the scarcity of benchmark resources in this field. Inspired by the UrbanWalk, NIG4VI utilizes the open-sourced CARLA Simulator (Dosovitskiy et al., 2017) to collect samples from a diverse range of scenarios. These scenarios span remote rural-like settings (e.g., Town01) and complex metropolitan areas (e.g., Town10), and encompass various weather conditions, such as foggy and sunny weather. Pedestrian trajectories are generated using A* route planning algorithm, with precise geospatial coordinates, orientation, frontal-view images, and semantic segmentation images being recorded at each step. NIG4VI offers two main advantages: (1) its use of a realistic coordinate system facilitates easier transfer to real-world GPS applications, and (2) it enables the cost-effective generation of accurate and extensive data. Table 1 demonstrates NIG4VI’s advantages compared to other datasets.
4.1 Dataset Construction
@@
Each question’s input includes the user’s current location/rotation, the next step’s location/orientation, and visual scene data. The complete prompt template used to structure these inputs is detailed in Table 6 of Appendix E. The synthesis of the output is a multi-stage process involving both advanced reasoning models and human annotation, as illustrated in Figure 3. Initially, several leading VLMs, specifically GPT-4o, Claude-3.5, and Gemini-2, generate predictions. Following a modality bridging approach, similar to that employed in Vision-R1 (Huang et al., 2025), these outputs are processed through DeepSeek-R1 to enhance blindness-oriented spatial guidance and navigability. The specific prompt guiding DeepSeek-R1 in this refinement task is detailed in Table 7.
Crucially, all instructions undergo rigorous human verification. This task is carried out by two annotators, both proficient in English and holding at least an undergraduate-level education, following a similar practice in (Zhao et al., 2024). The verification involves a two-stage process: first, one annotator performs initial content adjustments, adhering to task requirements. Subsequently, the second annotator reviews and verifies this work. Throughout this entire process, both annotators focus on ensuring: (1) elimination of visual references (e.g., color-based descriptors), (2) validation of non-visual landmarks, and (3) confirmation of metric precision for mobility-critical parameters.
4.2 Dataset Statistics
@@
Table 2 details the statistics of the dataset, which comprises routes collected from six distinct towns within the CARLA simulator. Further details on route sampling in the CARLA simulator are in Appendix D. On average, each town contributes approximately 26.2 navigation routes. The average Euclidean distance between the start and end points of these routes is 113.51 units, with an average of 353.8 steps required for completion. After deduplication, the dataset yielded an average of 2,222.7 step-level (image, question) samples per town. It is partitioned into a training set of 1,500 samples from Town01 and a test set. The test set comprises the remaining 613 intra-town samples from Town01, along with all inter-town samples from Town02 (2,579), Town03 (2,260), Town04 (2,316), Town05 (1,935), and Town10 (2,133). Each data sample is available in two versions: ‘with pre-calculation’ and ‘without pre-calculation’. The ‘without pre-calculation’ version requires the VLM to independently calculate navigational parameters (e.g., distance, direction), presenting a greater challenge in guidance generation. Conversely, the ‘with pre-calculation’ version provides the VLM with basic mathematical movement information. The VLM must then validate this data and assess the surroundings to generate the final navigation instruction.
Town Routes Avg dist. Avg steps # Samples
Town01 25 111.41 401 1,500 / 613
Town02 26 99.38 327 2,579
Town03 25 128.23 409 2,260
Town04 26 131.49 337 2,316
Town05 25 107.81 288 1,935
Town10 30 102.74 361 2,133
Avg. 26.2 113.51 353.8 2,222.7
Table 2: Statistics for sample routes. Dataset: 1,500 Town01 samples for training; the remaining 613 (Town01) and all other town samples for testing. ‘Avg dist.’: average Euclidean distance (route start to end). ‘Avg steps’: average steps per route. ‘# Samples’: deduplicated step-level (image, question) samples per town.
@@
Zero- (LaF-GRPO) Qwen-VL-3B 3.292 0.230 0.248 0.230 3.972 0.255 0.259 0.244
@@
Zero- (LaF-GRPO) Qwen-VL-3B 3.798 0.249 0.280 0.261 4.584 0.271 0.288 0.274
@@
Table 3: Evaluation results on the NIG4VI dataset across Intra-town and Inter-town subsets. Gray cells indicate results pertaining to Qwen2.5-VL models. Blue values highlight the best performing Qwen2.5-VL model within the Zero-Shot and Zero-(LaF-GRPO) categories. Bold values represent the highest score for each metric under a specific setting (with / without pre-calculation), while underlined values indicate the second-highest score.
5 Experimental Settings
Pre-Cal. Reward Types Intra-town ($N$ = 613) Inter-town ($N$ = 11,223)
@@ 5 Experimental Settings
Table 4: Ablation study results for the Qwen2.5-VL-3B model on the NIG4VI dataset with different reward functions. Bold values represent the highest score for each metric under its specific pre-calculation condition.
Pre-Cal. Model Intra-town ($N$ = 613) Inter-town ($N$ = 11,223)
@@
Table 5: Ablation study results for the Qwen2.5-VL-7B model on the NIG4VI dataset with varying training sample sizes. Bold values represent the highest score for each metric under its specific pre-calculation condition.
Dataset.
Experiments utilized the NIG4VI dataset, comprising Intra-town ($N=613$) and Inter-town ($N=11,223$) test subsets, under ’with/without pre-calculation’ conditions.
Models.
[…]
Evaluation Metrics.
Following previous studies in NIG (Huang et al., 2022; Fan et al., 2024; Kong et al., 2024), model performance was evaluated using a suite of widely adopted metrics: BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), METEOR (Banerjee and Lavie, 2005), and SPICE (Anderson et al., 2016). For each of these metrics, higher scores denote superior performance.
Baselines.
We compare LaF-GRPO against two primary baseline methods: (1) Zero-shot: This involves models directly on NIG4VI without prior task-specific fine-tuning. (2) Supervised Fine-tuning (SFT): Models are fine-tuned to generate instructions from the input. Furthermore, we implement two variants of LaF-GRPO to understand its different operational modes: (a) Zero-(LaF-GRPO): LaF-GRPO is applied directly to the base model without SFT. (b) SFT+(LaF-GRPO): LaF-GRPO is applied to models that have first undergone SFT. Both LaF-GRPO variants utilize the proposed LLM-as-Follower reward mechanism.
Implementation Deatails.
LaF-GRPO training utilized a single NVIDIA H20 GPU (96 GB of memory). This hardware supports loading an 8B-param LLM (LLaMA-3-8B) and a 3B/7B-param Qwen2.5-VL model for LoRA (Hu et al., 2022) fine-tuning. The reward weights were configured as $(w_{dir},w_{dist},w_{alert})=(0.4,0.4,0.2)$ based on analysis of navigation failure factors, prioritizing spatial parameters over contextual alerts. Training on $3k$ samples took approximately $15$ hours, with the key hyperparameter group size $G$ set to $8$.
6 Results and Discussions
6.1 Main Results
@@
We present Table 3, which summarizes model performance on NIG4VI, categorized by pre-calculation and training paradigms, and evaluated on intra-town and inter-town subsets. Comparing LaF-GRPO with the baselines reveals: (1) Zero-Shot vs. Zero-(LaF-GRPO): Zero-(LaF-GRPO) significantly enhances the Zero-Shot performance of VLMs, validating the effectiveness of LaF-GRPO. While the Zero-(LaF-GRPO) results suggest that increased model size (from 3B to 7B) does not necessarily guarantee improved performance across all metrics, it is noteworthy that for METEOR evaluations, specifically in intra-town scenarios, the 7B model achieved the highest scores (i.e., $0.256$ and $0.281$). This outcome may be attributable to the use of METEOR as a text generation reward during training and to the potentially more refined tuning applied to the 7B models. (2) SFT & SFT+(LaF-GRPO): SFT and SFT+(LaF-GRPO) yield significantly superior performance compared to Zero-Shot and Zero-(LaF-GRPO) models across all metrics and subsets, affirming the efficacy of fine-tuning. The SFT+(LaF-GRPO) approach further enhances performance beyond SFT. Moreover, under the SFT+(SFT-GRPO) paradigm, Qwen-VL-3B consistently achieves the highest BLEU and ROUGE scores, while Qwen-VL-7B excels in METEOR and SPICE. This performance pattern is observed for both intra-town and inter-town subsets and holds true regardless of pre-calculation. This may be attributable to 7B models demonstrating enhanced linguistic diversity in their outputs relative to 3B models. (3) Additional Observations: Scores are generally higher with pre-calculation than without, likely because it reduces the mathematical computation difficulty for the models. While both intra-town and inter-town results demonstrate consistent trends, intra-town evaluations typically yield higher scores. This is understandable as the intra-town test set comprises samples from the same town environments as the training set, leading to a closer data distribution.
6.2 Ablation Study
Reward Types
@@ 6.2 Ablation Study
Table 4 presents an ablation study investigating the impact of different reward types during SFT+(LaF-GRPO) training with the Qwen-VL-3B model. LaF-GRPO, incorporating the LLM-as-Follower reward, consistently achieves the highest BLEU, ROUGE, and METEOR scores. This trend holds true across both intra-town and inter-town evaluations, with or without pre-calculation. This underscores the significant benefit of the LLM-as-Follower reward signal for the NIG4VI task.
LaF-GRPO vs. Standard GRPO
We conducted an additional experiment on the inter-town subset. It reveals that LaF-GRPO-trained models demonstrate superior navigational accuracy ($68.1\%$ vs. $67.3\%$ ) and their instructions were more frequently selected by GPT-4o for helpfulness and clarity ($58.3\%$ vs. $41.7\%$ of cases). We identify two primary advantages of LaF-GRPO over standard GRPO, which utilizes only format and text generation rewards: (1) Navigational Accuracy: LaF-GRPO provides more precise movement and orientation accuracy. (2) Instruction Clarity: The inclusion of an action interpreter requirement encourages VLMs to produce instructions comprehensible to the follower and thus clearer and more structured. Details are provided in Appendix F.
Training Sample Sizes
@@
Table 5 presents an ablation study on the Qwen2.5-VL-7B model trained with SFT+(LaF-GRPO), illustrating the effect of varying training sample sizes ($1k$, $2k$, and $3k$). For the two metrics METEOR and SPICE, which are often considered more comprehensive in text generation, increasing the volume of training data generally leads to enhanced performance. Across the majority of evaluated conditions (with/without pre-calculation), scaling up to $3k$ samples typically yields the optimal or near-optimal scores. Nevertheless, for the 7B models in the inter-town setting, training with $2k$ samples also achieves comparable METEOR scores (i.e., $0.535$ and $0.541$), indicating training data efficiency at this sample size.
6.3 Case Study
@@
Figure 4 provides a qualitative comparison of our SFT+(LaF-GRPO) method against the SFT baseline. Notably, SFT+(LaF-GRPO) generates instructions with greater linguistic variety and more intuitive directional cues. For instance, in Step 2, SFT+(LaF-GRPO) employs an o’clock direction (“Turn slightly right toward 1 o’clock") and a relatable distance ("two small steps"), contrasting with SFT’s numerical bearing ("150 degrees"). This approach can yield guidance that is more naturally understood by VI users. Furthermore, SFT+(LaF-GRPO), leveraging its internal reasoning process (i.e. the <think>...</think> blocks), frequently incorporates more environmental details and safety considerations. For example, its instruction for Step 4 ("Step forward 0.5 meters; …use your cane near the left railing and listen for traffic") also emphasizes immediate safety interactions. Full instruction texts for Zero-(LaF-GRPO), SFT, and SFT+(LaF-GRPO) are available in Appendix F.

### v2

@@ 4 Benchmark: NIG4VI
R2R (DBLP:conf/cvpr/AndersonWTB0S0G18) High 21k
REVERIE (DBLP:conf/cvpr/QiW0WWSH20) High 10k / 6k
UrbanWalk (DBLP:conf/eccv/HuangSZBBO22) Detailed 2.6k
merchant2024generatingcontextuallyrelevantnavigationinstructions Detailed 48
VIALM (zhao2024vialmsurveybenchmarkvisually) Detailed 200
WAD (yuan2025walkvlmaidvisuallyimpairedpeople) Detailed 12k / 120k
@@
Table 1: Comparison of NIG4VI with existing benchmarks.
@@
We introduce the NIG4VI benchmark to address the scarcity of benchmark resources in this field. Inspired by UrbanWalk, NIG4VI utilizes the open-sourced CARLA Simulator (Dosovitskiy17) to collect samples from diverse scenarios, including varied environments and weather conditions. Pedestrian trajectories are generated using A* algorithm, with precise geospatial coordinates, orientation, frontal-view images, and semantic segmentation images recorded at each step. NIG4VI offers two advantages: (1) its use of a realistic coordinate system facilitates easier transfer to real-world GPS applications, and (2) it enables the cost-effective generation of accurate and extensive data. Table 1 details NIG4VI’s advantages over other datasets.
@@
Dataset Construction. Each question’s input includes the user’s current location/rotation, the next step’s location/orientation, and visual scene data. This information is then structured within a prompt template that includes a detailed task description. The synthesis of the output is a multi-stage process involving both advanced reasoning models and human annotation, as illustrated in Figure 3. Initially, several leading VLMs, specifically GPT-4o, Claude-3.5, and Gemini-2, generate predictions. Following a modality bridging approach, similar to that employed in Vision-R1 (huang2025visionr1incentivizingreasoningcapability), these outputs are processed through DeepSeek-R1 to enhance blindness-oriented spatial guidance and navigability. Crucially, all instructions undergo rigorous human verification. This task is carried out by two annotators, both proficient in English and holding at least an undergraduate-level education, following a similar practice in (zhao2024vialmsurveybenchmarkvisually). The verification involves a two-stage process: first, one annotator performs initial content adjustments, adhering to task requirements. Subsequently, the second annotator independently reviews and verifies this work. Throughout this process, the primary verification criteria include: (1) elimination of visual references (e.g., color-based descriptors), (2) validation of non-visual landmarks, and (3) confirmation of metric precision for mobility-critical parameters.
Figure 3: Dataset instances are first generated using Vision-R1’s modality bridging method (huang2025visionr1incentivizingreasoningcapability) and then reviewed and refined by human annotators.
@@
Dataset Statistics. On average, each town contributes approximately 26.2 navigation routes. The average Euclidean distance between the start and end points of these routes is 113.51 units, with an average of 353.8 steps. After deduplication, the dataset yielded an average of 2,222.7 step-level (image, question) samples per town. It is partitioned into a training set of 1,500 samples from Town01 and a test set. The test set comprises the remaining 613 intra-town samples from Town01, along with all inter-town samples from Town02 (2,579), Town03 (2,260), Town04 (2,316), Town05 (1,935), and Town10 (2,133). Each data sample is available in two versions: ‘with pre-calculation’ and ‘without pre-calculation’. The ‘without pre-calculation’ version requires the VLM to independently calculate navigational parameters (e.g., distance, direction), presenting a greater challenge in guidance generation. Conversely, the ‘with pre-calculation’ version provides the VLM with basic mathematical movement information. The VLM must validate this data and assess the surroundings to generate the navigation instruction.
5 Experimental Settings
Dataset. Experiments utilized the NIG4VI dataset, with Intra-town ($N=613$) and Inter-town ($N=11,223$) test subsets, under ‘with/without pre-calculation’ conditions.
Models. Diverse VLMs were evaluated, falling into two main groups: (1) remote models, such as GPT-4o (openai2024gpt4ocard), Claude-3-5-sonnet-20240620 (Anthropic2024bClaudeSonnet), and Gemini-2.0-flash-thinking-exp-01-21 (GoogleDeepMind2024bGeminiIntroMisc); and (2) smaller, locally runnable VLMs, including DeepSeek-VL-7B (lu2024deepseekvl), LLaVA-v1.6-Mistral-7B (DBLP:conf/cvpr/LiuLLL24), MiniCPM-o-2.6-8B (yao2024minicpm), Intern-VL-2.5-8B (chen2024internvl), and Qwen2.5-VL-3B/7B (Qwen2.5-VL).
Baselines. We compare LaF-GRPO against two primary baseline methods: (1) Zero-shot: Models are applied directly to NIG4VI without prior fine-tuning. (2) Supervised Fine-tuning (SFT): Models are fine-tuned to generate instructions. We implement two variants of LaF-GRPO: (a) Zero-(LaF-GRPO): LaF-GRPO is applied directly to the base model without SFT. (b) SFT+(LaF-GRPO): LaF-GRPO is applied to models that have first undergone SFT.
Evaluation Metrics. Following previous studies in NIG (DBLP:conf/eccv/HuangSZBBO22; DBLP:conf/eccv/FanLWY24; DBLP:conf/eccv/KongCWSHYL24), model performance was evaluated using a suite of widely adopted metrics: BLEU (DBLP:conf/acl/PapineniRWZ02), ROUGE (lin-2004-rouge), METEOR (banerjee-lavie-2005-meteor), and SPICE (anderson2016spicesemanticpropositionalimage). Human studies and LLM-as-Judge evaluations were conducted to assess navigational accuracy and user preference for instruction clarity and helpfulness.
Implementation Details. LaF-GRPO training utilized a single NVIDIA H20 GPU (96 GB of memory). This hardware supports loading an 8B-param LLM (LLaMA-3-8B) and a 3B/7B-param Qwen2.5-VL model for LoRA (hu2022lora) fine-tuning. The reward weights were configured as $(w_{dir},w_{dist},w_{alert})=(0.4,0.4,0.2)$ based on analysis of navigation failure factors, prioritizing spatial parameters over contextual alerts. Training on $3k$ samples took about $15$ hours, with the hyperparameter group size $G$ set to $8$.
6 Results and Discussions
@@
Zero-
(LaF-GRPO)
Qwen-VL-3B 3.292 0.230 0.248 0.230 3.972 0.255 0.259 0.244
@@
Zero-
(LaF-GRPO)
Qwen-VL-3B 3.798 0.249 0.280 0.261 4.584 0.271 0.288 0.274
@@
Table 2: Evaluation results on NIG4VI across Intra-town and Inter-town subsets. Bold values represent the highest score for each metric under a specific setting (with / without pre-calculation), while underlined values indicate the second-highest score. Boxed values highlight the best performing Qwen2.5-VL model within the Zero-Shot and Zero-(LaF-GRPO) categories.
Pre-Cal.
Reward Types Intra-town ($N$ = 613) Inter-town ($N$ = 11,223)
@@
Table 3: Ablation study results for the Qwen2.5-VL-3B model on the NIG4VI dataset with different reward functions.
Pre-Cal.
Model Intra-town ($N$ = 613) Inter-town ($N$ = 11,223)
@@
Table 4: Ablation study results for the Qwen2.5-VL-7B model on the NIG4VI dataset with varying training sample sizes.
Main Results
@@
Table 2 summarizes model performance on NIG4VI, categorized by pre-calculation and training paradigms, and evaluated on intra-town and inter-town subsets. Comparing LaF-GRPO with the baselines reveals: (1) Zero-Shot vs. Zero-(LaF-GRPO): Zero-(LaF-GRPO) significantly enhances the Zero-Shot performance of VLMs, validating the effectiveness of LaF-GRPO. While the Zero-(LaF-GRPO) results suggest that increased model size (from 3B to 7B) does not necessarily guarantee improved performance across all metrics, it is noteworthy that for METEOR evaluations, specifically in intra-town scenarios, the 7B model achieved the highest scores (i.e., $0.256$ and $0.281$). This outcome may be attributable to the use of METEOR as a text generation reward during training and to the potentially more refined tuning applied to the 7B models. (2) SFT & SFT+(LaF-GRPO): SFT and SFT+(LaF-GRPO) yield significantly superior performance compared to Zero-Shot and Zero-(LaF-GRPO) models across all metrics and subsets, affirming the efficacy of fine-tuning. The SFT+(LaF-GRPO) approach further enhances performance beyond SFT. Moreover, under the SFT+(LaF-GRPO) paradigm, Qwen-VL-3B consistently achieves the highest BLEU and ROUGE scores, while Qwen-VL-7B excels in METEOR and SPICE. This performance pattern is observed for both intra-town and inter-town subsets and holds true regardless of pre-calculation. This may be attributable to 7B models demonstrating enhanced linguistic diversity in their outputs relative to 3B models. (3) Additional Observations: Scores improved with pre-calculation, which reduced computational difficulty, whereas the more challenging ‘w/o pre-calculation’ setting caused more failures. Intra-town scores were higher than inter-town, likely due to a closer data distribution with the training set. We also observed that LaF-GRPO generated significantly more concise, user-friendly instructions (e.g., LaF-GRPO-7B:  34.1 tokens vs. GPT-4o:  117.9 tokens).
Ablation Studies
@@
Reward Types. Table 3 presents an ablation study investigating the impact of different reward types during SFT+(LaF-GRPO) training with the Qwen-VL-3B model. LaF-GRPO, incorporating the LLM-as-Follower reward, consistently achieves the highest BLEU, ROUGE, and METEOR scores. This trend holds true across both intra-town and inter-town evaluations, with or without pre-calculation. This underscores the significant benefit of the LaF reward.
@@
Training Sample Sizes. Table 4 presents an ablation study on the Qwen2.5-VL-7B model trained with SFT+(LaF-GRPO), illustrating the effect of varying training sample sizes ($1k$, $2k$, and $3k$). For comprehensive metrics such as METEOR and SPICE, performance generally scales with the volume of training data. Across the majority of evaluated conditions, scaling up to $3k$ samples typically yields the optimal or near-optimal scores. Nevertheless, training with $2k$ samples also achieves comparable METEOR scores, indicating training data efficiency at this sample size.
LaF-GRPO vs. Standard GRPO. We conducted LLM-as-Judge and human evaluations on the inter-town subset. The LLM-as-Judge evaluation with GPT-4o showed LaF-GRPO’s superior navigational accuracy (68.1% vs. 67.3%) and a greater preference rate for instruction clarity and helpfulness (58.3% vs. 41.7%). These findings were corroborated by a human preference study on 100 instruction pairs, evaluated by two trained graduate students acting as VI user simulators. To isolate instruction quality, the evaluators were instructed to navigate relying solely on text and ignoring visual cues. Their assessments achieved substantial inter-rater agreement (Cohen’s $\kappa$= 0.83), and the results confirmed a strong preference for LaF-GRPO instructions (76% vs. 24%) and its higher navigational accuracy (79.0% vs. 77.5%). We identify two main advantages of LaF-GRPO over standard GRPO, which utilizes only format and text generation rewards: (1) Navigational Accuracy: LaF-GRPO provides more precise movement and orientation guidance. (2) Instruction Clarity: The action interpreter component encourages VLMs to generate instructions that are clearer, more structured, and more comprehensible.
Case Study
@@
Figure 4 provides a qualitative comparison of SFT+(LaF-GRPO) against the SFT baseline. Notably, SFT+(LaF-GRPO) generates instructions with greater linguistic variety and more intuitive directional cues. For instance, in Step 2, SFT+(LaF-GRPO) employs an o’clock direction (“Turn slightly right toward 1 o’clock”) and a relatable distance (”two small steps”), contrasting with SFT’s numerical bearing (”150 degrees”). It can yield guidance that is more naturally understood by VI users. Furthermore, SFT+(LaF-GRPO), leveraging its internal reasoning process (i.e. the <think>...</think> blocks), frequently incorporates more environmental details and safety considerations. For example, its instruction for Step 4 (”Step forward 0.5 meters; …use your cane near the left … and listen for traffic”) also emphasizes immediate safety interactions.
Limitations and Future Work
This study’s reliance on simulation and proxy users, while necessary for large-scale, controlled, and safe initial testing, introduces limitations. Future work could explore real-world dataset collection and direct engagement with VI users.